OpenStudy (anonymous):

Circumference = 2πr Circumference = 30 All I know so far is , 30 = 2πr

OpenStudy (anonymous):

divide both sides by 2pi to solve for r. r=30/2pi=4.77

OpenStudy (anonymous):

so you would solve for r. r=30/2pi r=4.774648293

OpenStudy (anonymous):

I have to show my work though , that's the problem.. I don't know how. Is this how.. (hold on let me draw what im randomly doing)

OpenStudy (anonymous):

 that's all I know.